Dawson Toys, Ltd., produces a toy called the Maze. The company has recently established a standard cost system to help control costs and has established the following standards for the Maze toy: Direct materials: 7 microns per toy at $0.30 per micron Direct labor: 1.4 hours per toy at $6.50 per hour During July, the company produced 5,400 Maze toys. The toy's production data for the month are as follows: Direct materials: 79,000 microns were purchased at a cost of $0.28 per micron. 31,750 of these microns were still in inventory at the end of the month. Direct labor: 7,860 direct labor-hours were worked at a cost of $55,020. Required: 1. Compute the following variances for July: (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values. Do not round intermediate calculations. Round final answer to the nearest whole dollar amount.) a. The materials price and quantity variances. b. The labor rate and efficiency variances.
Answer #1
Standard Actual
Particulars Qty/ hours Rate amount Qty/ hours Rate amount
Materials            37,800.00                      0.30            11,340.00            79,000.00                      0.28            22,120.00
Labour              7,560.00                  6.5000            49,140.00              7,860.00                      7.00            55,020.00
Actual output              5,400.00
Materials reqd (5400*7)            37,800.00
Labour hrs reqd (5400*1.4)              7,560.00
AQ Used = 79000 - 31750            47,250.00
DMPV = (SP-AP)*AQ purchased
DMPV = (.30 - .28)79000
DMPV = 1,580 F
DMQV= (SQ-AQ)SP
DMQV= (37800 - 47250).30
DMQV= 2,835 U
DLRV= (SR-AR)AH
DLRV= (6.5 - 7)7860
DLRV= 3,930 U
DLEV = (SH-AH)SR
DLEV = (7560 - 7860)6.5
DLEV = 1,950 U